Please help solve the equation with steps thank you
3x^3+19x^2-72x=0

Answer by jorel555(1290)

3x³+19x²-72x=0
x(3x²+19x-72)=0
x(3x-8)(x+9)=0
x=0,8/3, or -9. ☺☺☺☺